OpenStudy (abbycross167):

Graph the following equations using the point-slope method or intercepts method. (1 pt) Show the steps of the procedure as outlined in the lesson as you graph the lines. (1 pt) Draw your y-intercept in yellow, your other points in red and your line in blue. x + 3y= -6

OpenStudy (abbycross167):

@Directrix can you try to help me with this one? I have no clue how to do this one

Directrix (directrix):

Use the intercepts method: The x intercept lies on the x axis. The y coordinate of all points on the x axis is zero. Let y = 0 and solve for x. x + 3y= -6 x + 3*0 = -6 @abbycross167 Solve for x.

OpenStudy (abbycross167):

x+0=-6 x= -6 that is as far as I got doing it by myself

OpenStudy (abbycross167):

@Directrix

Directrix (directrix):

The y intercept lies on the y axis. The x coordinate of all points on the y axis is zero. Let x = 0 and solve for y. 0 + 3y= -6 Solve for y.

OpenStudy (monkeyagf):

pls help me when you guys are done with this question

OpenStudy (abbycross167):

-6/3= -2?

Directrix (directrix):

Be careful not to lose the variable when you are solving equations. 0 + 3y= -6 3y = -6 y = -2 You have the -2 but lost whether it is the value of y or x.

Directrix (directrix):

Plot the points (6,0) and (0,-2). Then draw the line connecting them.

OpenStudy (abbycross167):



OpenStudy (abbycross167):

like that?

Directrix (directrix):